### v1

# The Geometric Reasoner: Manifold-Informed Latent Foresight Search for Long-Context Reasoning

Abstract
[…]
Test-Time Compute, Latent-Manifold Search, Value-Based Look-Ahead, Geometric Regularization, Inference-Time Scaling, Long-Context Reasoning, Large Language Models
[…]
1 Introduction
Scaling test-time compute reliably enhances large language model (LLM) performance, unlocking reasoning capabilities previously out of reach (Xu et al., 2025; Muennighoff et al., 2025; Zhang et al., 2025b; Wu et al., 2025). Yet turning extra compute into diverse exploration without incurring prohibitive memory costs or wasting budget on redundant, correlated trajectories remains difficult. The quadratic complexity of attention and the footprint of KV caching (Vaswani et al., 2017; Beltagy et al., 2020; Dao et al., 2022; Dao, 2023) bottleneck practical deployment, rendering naive long-horizon exploration intractable.
Existing approaches make progress on this tension, yet critical gaps persist. Reinforcement Learning (RL) and preference optimization amortize long-horizon control into model weights (Schulman et al., 2017; Christiano et al., 2017; Shao et al., 2024; Zheng et al., 2025). While effective for single-sample accuracy, these methods demand substantial training compute and can collapse the trajectory distribution (Yue et al., 2025; Srivastava and Aggarwal, 2025). Sampling-based inference trades compute for robust coverage without retraining (Wang et al., 2022; Karan and Du, 2025; Faria et al., 2024), yet returns diminish rapidly as additional samples often remain correlated and fail to expand the set of distinct, valid trajectories (Tang et al., 2025; Wu et al., 2025). These limitations motivate a training-free inference-time search procedure that allocates compute dynamically under strict memory bounds.
Figure 1: TGR-Latent consistently outperforms baselines on Qwen3-8B. Manifold-informed latent foresight search steering converts modest inference-time compute into robust coverage without weight updates.
To address this, we introduce The Geometric Reasoner (TGR), a training-free framework for manifold-informed latent foresight search. TGR formulates reasoning as path search on a latent manifold but avoids rigid constraints that become brittle in high dimensions (Yallup et al., 2025; Betancourt, 2017; Goyal and Shetty, 2019). Instead, it ranks candidate reasoning chunks via a soft geometric score combining lightweight look-ahead utility, a bumpiness penalty for smooth transitions, and a repulsive diversity force that discourages similarity. Chunk-wise KV cache resets keep memory linear in chunk length, enabling scalable long-context exploration.
This chunk-boundary state interface relates to markovian control approaches that learn explicit state transitions, like Delethink (Aghajohari et al., 2025). Rather than relying on learned state transitions or external verifiers, TGR enables training-free inference-time search, allocating compute to achieve controllable exploration and robust coverage under matched budgets.
Our contributions are:
  • Manifold-Informed Latent Foresight Search. We propose TGR, a training-free inference-time framework that scores chunk-level latent anchors with a soft geometric objective, enabling controllable exploration without parameter updates.
[…]
  • Scaling Limits of Hard Geometry. We demonstrate that hard geometric filtering suffers from vanishing feasibility in high-dimensional latent spaces, where the acceptance rate decays exponentially as dimension increases, establishing soft geometric regularization as a scalable alternative.
2 Related Work
Figure 2: Overview of reasoning frameworks. Unlike (a) test-time sampling, exploring trajectories without explicit structure, or (b) reinforcement learning, which internalizes preferences through costly training, (c) TGR introduces a training-free inference-time search over the latent manifold. It selects optimal chunk-level anchors via a soft geometric score combining foresight, bumpiness, and uniformity, then injects them into generation with KV cache resets to maintain linear memory cost.
2.1 Test-time Reasoning
Prompt engineering steers model behavior via input formatting and exemplars, spanning in-context learning (Brown et al., 2020), Chain-of-Thought decomposition (Wei et al., 2022; Kojima et al., 2022), and program-aided methods such as PAL (Gao et al., 2023). These approaches improve controllability without additional training but remain sensitive to prompt design and operate via token-level heuristics.
Sampling-based reasoning includes local decoding controls (temperature, top-$k$, top-$p$), candidate pooling with Best-of-$N$ selection (Wan et al., 2025), and aggregation methods such as Self-Consistency (Wang et al., 2022). Aggressive test-time sampling shows that additional inference compute can elicit latent reasoning without retraining (Karan and Du, 2025). Approaches such as Tree of Thoughts (ToT) (Yao et al., 2023) and Graph of Thoughts (GoT) (Besta et al., 2024) exemplify structured inference, which introduces explicit search in token space and trades computation for robustness. However, these methods search over discrete continuations and rely on token-space proxies or external scoring, rendering long-horizon search brittle under limited budgets.
2.2 Reinforcement Learning and Preference Optimization
Amortizing long-horizon control into model parameters via reinforcement learning and preference optimization has become a dominant paradigm (Christiano et al., 2017; Zhang et al., 2025a; Wiering and Van Otterlo, 2012). Beyond PPO (Schulman et al., 2017) and DPO (Rafailov et al., 2023), recent methods including GRPO (Shao et al., 2024) and GSPO (Zheng et al., 2025) scale reasoning by reducing reliance on a learned critic. Delethink (Aghajohari et al., 2025) learns discrete state transitions, demonstrating the utility of explicit Markovian states for long-horizon control. While effective, reward-and-preference-driven training exhibits failure modes at reduced diversity and overconfident generations (Yue et al., 2025; Srivastava and Aggarwal, 2025). SimKO (Peng et al., 2025) mitigates mode collapse and length bias, yet these training-centric approaches remain computationally intensive.
2.3 Geometric Perspectives on Latent Modeling
Geometric Deep Learning interprets learned representations as structured spaces (Bronstein et al., 2017; Higgins et al., 2017). Recent works frame attention as a geometric object governing representation transport along a curved manifold, treating layers as discrete geodesic steps shaped by training (Aggarwal et al., 2025; Di Sipio et al., 2025; Zhu et al., 2025). Complementing this view, LLM decoding is interpreted as tracing trajectories on a low-dimensional manifold, where coherent generations follow smooth paths and hallucinations appear as abrupt jumps (Zhang and Dong, 2025). These insights motivate geometric inductive biases at training time: manifold priors (Kang et al., 2025) and architectural constraints such as mHC (Xie et al., 2025), which applies Sinkhorn–Knopp normalization (Sinkhorn and Knopp, 1967) to confine signal propagation to a stable low-dimensional structure. Soft enforcement of geometry is preferred over hard feasibility filtering, which becomes impractical in high dimensions due to concentration of measure (Yallup et al., 2025; Betancourt, 2017; Goyal and Shetty, 2019). TGR operationalizes this insight at inference time, translating geometric priors into differentiable soft penalties without architectural changes.
Figure 2 contrasts TGR with test-time reasoning and RL methods, highlighting its training-free, manifold-informed latent foresight search.
3 Method
TGR is a training-free geometric reasoner that operates in latent space. It keeps the backbone LLM frozen, segments reasoning into chunks, and performs search at chunk boundaries within a continuous anchor space. At each boundary, TGR samples candidate anchors, scores them using computationally efficient rollouts and soft geometric penalties, and selects the highest-scoring anchor to steer the subsequent chunk. Note that in this work, TGR denotes the latent-space variant TGR-Latent, unless explicitly stated otherwise.
3.1 Latent Space State Representation
Given a query $Q$, TGR generates output as a sequence of $L$ chunks $\tau_{1:L}=(\tau_{1},\ldots,\tau_{L})$, each of maximum length $S$. Let $c_{t-1}$ denote the token context preceding chunk $t$. To bound memory during long-horizon decoding, we employ chunk-wise KV cache resets: at each boundary, we rebuild the prefix cache from $c_{t-1}$, discarding earlier history while transmitting long-range state via latent anchors (Appendix LABEL:app:impl). Concretely, $c_{t-1}$ comprises the query and recent token suffix, left-truncated to maximum length $S$.
3.1.1 State anchors
We represent chunk-level reasoning states as unit-norm latent anchors on $\mathcal{Z}\triangleq\{x\in\mathbb{R}^{d_{z}}:\lVert x\rVert_{2}=1\}$. A state anchor $z_{t}\in\mathcal{Z}$ is extracted upon completing chunk $t$, serving as a compact chunk-boundary summary. Unlike Delethink (Aghajohari et al., 2025), which learns discrete state transitions, TGR extracts anchors directly from hidden states and performs score-guided anchor selection at inference time. This provides a summary for control under bounded KV cache constraints. At the next boundary, we sample $K$ candidate anchors $\{\hat{z}_{t}^{(j)}\}_{j=1}^{K}\subset\mathcal{Z}$ centered at $z_{t-1}$, evaluate each via lightweight rollouts, and select the highest-scoring candidate to steer the next chunk.
This design addresses two inference constraints: chunk-wise KV resets require explicit state summaries for long-horizon coherence, and effective compute allocation demands a stateful interface rather than unstructured sampling.
3.1.2 State extraction
Upon generating $\tau_{t}$, we append a fixed end-of-chunk delimiter token <EOC> and project its top-layer hidden state $h_{\mathrm{eoc}}(\tau_{t})\in\mathbb{R}^{d_{h}}$ onto the unit sphere , equivalently, the last hidden state of $\tau_{t}\oplus\texttt{\textless EOC\textgreater}$:
$z_{t}\triangleq\frac{Wh_{\mathrm{eoc}}(\tau_{t})}{\lVert Wh_{\mathrm{eoc}}(\tau_{t})\rVert_{2}}\in\mathcal{Z},$ (1)
where $W\in\mathbb{R}^{d_{z}\times d_{h}}$ is a fixed projection matrix. This yields a compact, model-internal summary facilitating chunk-to-chunk state transfer despite context truncation.
3.2 Foresight Search
We generate candidate anchors by sampling from the local neighborhood of the current state anchor via tangent-space perturbation followed by re-normalization:
$\epsilon$ $\sim\mathcal{N}(0,I),$ (2)
$v$ $\leftarrow\epsilon-(z^{\top}\epsilon)z,$
$\tilde{a}$ $\leftarrow z+\sigma v,$
$a$ $\leftarrow\frac{\tilde{a}}{\lVert\tilde{a}\rVert_{2}}\in\mathcal{Z},$
where $z$ is the center anchor, setting $z=z_{t-1}$, and $\sigma$ controls the exploration radius. Tangent-space perturbations preserve the local neighborhood structure on the unit sphere, avoiding radial drift associated with ambient space rescaling. We sample $K$ candidate anchors $\{\hat{z}_{t}^{(j)}\}_{j=1}^{K}$ around $z_{t-1}$ using Eq. (2). Each candidate is evaluated using the manifold-informed soft geometric score (Section 3.3), and the optimal anchor is selected to generate the subsequent chunk.
3.3 Manifold-Informed Regularization
We operationalize soft geometry as manifold-informed regularization that exploits local structure to guide search, leveraging tangent neighborhoods and curvature proxies. Unlike hard feasibility filtering, our approach translates geometric priors into differentiable soft penalties.
At each boundary, we score a candidate anchor $a$ given context $c$ and previous anchor $z$ by maximizing:
$\mathrm{Score}(a;c,z)$ $=V_{\mathrm{fore}}(a;c)$ (3)
$\quad-\lambda_{b}\,P_{\mathrm{bum}}(a;c)-\lambda_{u}\,P_{\mathrm{uni}}(a;z).$
Here, $\lambda_{b},\lambda_{u}\geq 0$ balance capability, coherence, and exploration.
3.3.1 Foresight value.
We estimate look-ahead value via a low-depth rollout of $s\ll S$ tokens under $p_{\theta}(\cdot\mid c,a)$, where $p_{\theta}$ denotes the frozen base model conditioned on anchor $a$ via residual injection:
$V_{\mathrm{fore}}(a;c)$ $\triangleq\mathbb{E}\left[\frac{1}{s}\sum_{i=1}^{s}\log p_{\theta}(\hat{\tau}_{i}\mid\hat{\tau}_{<i};c,a)\right],$ (4)
$\qquad\hat{\tau}_{1:s}\sim p_{\theta}(\cdot\mid c,a).$
We approximate this expectation with a single Monte Carlo rollout per candidate. In this setting, $V_{\mathrm{fore}}$ becomes a single-sample average log-likelihood proxy sufficient for ranking. This direct measurement of consistency in low-cost rollouts enables ranking without external reward models or trained verifiers.
3.3.2 Bumpiness penalty.
Let $\mathbf{g}_{1},\ldots,\mathbf{g}_{s}$ be the top-layer hidden states along the rollout. We penalize high-frequency variation in the latent trajectory via a discrete second-order difference:
$P_{\mathrm{bum}}(a;c)\triangleq\frac{1}{s-2}\sum_{i=2}^{s-1}\left\lVert\mathbf{g}_{i+1}-2\mathbf{g}_{i}+\mathbf{g}_{i-1}\right\rVert_{2}^{2}.$ (5)
This term discourages abrupt directional changes, functioning as an efficient proxy for cross-chunk coherence.
3.3.3 Uniformity regularizer.
Since $a,z\in\mathcal{Z}$ are normalized, their semantic similarity reduces to a dot product. We enforce uniformity through repulsive interactions via a hinge penalty with threshold $\delta\in[0,1)$, thereby promoting diverse exploration:
$P_{\mathrm{uni}}(a;z)\triangleq\max\{0,\,a^{\top}z-\delta\}.$ (6)
This penalizes minimal-drift updates only when similarity exceeds $\delta$ (default $\delta=0.2$, see Appendix LABEL:app:hyperparams).
3.4 Overall Process
Algorithm 1 details the TGR inference process. At each boundary, we sample $K$ candidate anchors, compute the manifold-informed score using efficient rollouts, select the highest-scoring anchor, generate the next chunk, and update the state anchor. Formally, letting $\mathcal{A}_{t}\triangleq\{\hat{z}_{t}^{(j)}\}_{j=1}^{K}$ be the candidate set at step $t$, we select:
$\hat{z}_{t}\triangleq\arg\max_{a\in\mathcal{A}_{t}}\mathrm{Score}(a;c_{t-1},z_{t-1}).$ (7)
Algorithm 1 TGR Reasoning Process
0:  Query $Q$, chunk limit $L$, chunk length $S$, candidate anchors $K$, rollout length $s$.
0:  Proposal radius $\sigma$, weights $\lambda_{b},\lambda_{u}$, threshold $\delta$.
0:  Frozen base model parameters $\theta$; fixed projection $W$ defined in Eq. 1; fixed injectors $\{A_{\ell},B_{\ell}\}$ defined in Eq. 8.
0:  Output trajectory $\tau_{1:T}$.
1:  Initialize context $c_{0}\leftarrow\mathrm{Window}(Q)$.
2:  Run frozen model on $c_{0}$, set $z_{0}\leftarrow\text{Normalize}(Wh_{\mathrm{eoc}}(c_{0}))$ via Eq. 1.
3:  for $t=1,\ldots,L$ do
4:  Rebuild prefix KV cache for $c_{t-1}$ via chunk-wise reset.
5:  for $j=1,\ldots,K$ do
6:   Sample candidate $\hat{z}_{t}^{(j)}\leftarrow\textsc{SampleAround}(z_{t-1},\sigma)$ using Eq. 2.
7:   Roll out $\hat{\tau}^{(j)}_{1:s}\sim p_{\theta}(\cdot\mid c_{t-1},\hat{z}_{t}^{(j)})$ for $s$ steps, applying residual injection via Eq. 8.
8:   Compute foresight value $V_{\mathrm{fore}}(\hat{z}_{t}^{(j)};c_{t-1})$ via Eq. 4.
9:   Compute bumpiness $P_{\mathrm{bum}}(\hat{z}_{t}^{(j)};c_{t-1})$ via Eq. 5.
10:   Compute uniformity $P_{\mathrm{uni}}(\hat{z}_{t}^{(j)};z_{t-1})$ via Eq. 6.
11:   Score $\mathcal{S}^{(j)}\leftarrow V_{\mathrm{fore}}-\lambda_{b}P_{\mathrm{bum}}-\lambda_{u}P_{\mathrm{uni}}$ using Eq. 3.
12:  end for
13:  Select $j^{\star}\leftarrow\arg\max_{j}\mathcal{S}^{(j)}$; set $\hat{z}_{t}\leftarrow\hat{z}_{t}^{(j^{\star})}$.
14:  Generate chunk $\tau_{t}\sim p_{\theta}(\cdot\mid c_{t-1},\hat{z}_{t})$, length $\leq S$, via injection as in Eq. 8.
15:  Extract new anchor $z_{t}\leftarrow\text{Normalize}(Wh_{\mathrm{eoc}}(\tau_{t}))$ via Eq. 1.
16:  Update context $c_{t}\leftarrow\mathrm{Window}(Q\oplus\tau_{t})$ resetting KV cache.
17:  if stopping criterion met then
18:   break
19:  end if
20:  end for
21:  return $\tau_{1:t}$
Anchor conditioning via residual injection
We condition on anchor $a_{t}\in\mathcal{Z}$ using a fixed low-rank residual injection at each transformer layer $\ell$:
$\mathbf{h}_{\ell}$ $\leftarrow\mathbf{h}_{\ell}+B_{\ell}A_{\ell}a_{t},$ (8)
$A_{\ell}$ $\in\mathbb{R}^{r\times d_{z}},\quad B_{\ell}\in\mathbb{R}^{d_{h}\times r},\quad r\ll d_{h}.$
This creates a lightweight control interface without parameter updates. We inject at every layer using identical interfaces for candidate evaluation and generation, with fixed random matrices of rank $r$ (default $r=8$) to maintain the training-free setting (Appendix LABEL:app:impl).
3.5 Vanishing Acceptance of Hard Geometric Constraints
A standard method for imposing latent geometric structure is hard feasibility filtering—accepting candidates only if they satisfy constraints such as geodesic-following, curvature caps, or angular thresholds.
Let $a$ denote a candidate sampled from proposal distribution $q(\cdot\mid z_{t-1})$ centered at $z_{t-1}$, and let $\mathcal{F}^{\mathrm{geo}}_{t}$ be the feasible set induced by a hard constraint, namely a curvature-bounded geodesic-following rule. The feasibility acceptance rate, i.e., pure accept/reject excluding likelihood ratios, is:
$\begin{split}\alpha\triangleq{}&\Pr_{a\sim q(\cdot\mid z_{t-1})}\!\big[a\in\mathcal{F}^{\mathrm{geo}}_{t}\big]\\ =&\mathbb{E}_{a\sim q(\cdot\mid z_{t-1})}\!\left[\mathbb{I}\{a\in\mathcal{F}^{\mathrm{geo}}_{t}\}\right].\end{split}$ (9)
The expected number of proposals required to obtain one feasible sample is $\mathbb{E}[N_{\mathrm{prop}}]=1/\alpha$. Maintaining a fixed number of feasible candidates, such as $K$ candidates per step, therefore requires inflating the proposal budget by a factor of roughly $1/\alpha$.
Crucially, $\alpha$ collapses rapidly in high-dimensional latent spaces due to concentration of measure; feasible regions often occupy an exponentially small fraction of the local neighborhood as dimensionality increases (Betancourt, 2017; Goyal and Shetty, 2019; Yallup et al., 2025). Table LABEL:tab:hard_acceptance illustrates that hard filtering becomes computationally prohibitive as dimension increases.
This motivates our use of soft geometry via manifold-informed regularization. Rather than rejecting candidates outside $\mathcal{F}^{\mathrm{geo}}_{t}$, TGR integrates geometric priors as differentiable penalties within the scoring objective. This approach preserves geometric bias while circumventing the prohibitive $1/\alpha$ cost penalty of hard feasibility filtering.
Table 1: Mathematical reasoning results on Qwen3-8B. TGR-Latent matches or exceeds the best baseline on AUC while consuming 18% fewer tokens, demonstrating that inference-time latent search achieves superior budget-efficiency without training updates.
Method AIME25 OlympiadBench OmniMath Tokens
@1 @32 @128 AUC @1 @32 @128 AUC @1 @32 @128 AUC Avg. ($\times 10^{3}$)
Base Model 18.7 26.5 28.4 24.0 22.1 30.2 32.6 27.9 31.4 45.8 48.9 41.2 2.6
PSampling 23.8 30.1 32.1 28.4 27.4 34.5 36.8 32.6 38.7 49.8 52.3 46.2 6.0
LCoT-GRPO 26.5 27.8 28.5 27.5 30.1 31.8 32.5 31.5 41.8 48.0 49.8 46.2 8.5
LCoT-SimKO 27.5 31.5 33.0 30.5 31.8 36.0 38.0 35.1 43.5 51.5 54.0 49.1 8.6
Dele-GRPO 27.4 35.5 37.8 33.2 31.5 39.0 41.8 37.2 43.1 54.2 57.1 50.9 8.4
Dele-SimKO 28.4 37.0 39.5 34.8 32.5 40.8 43.8 38.8 44.6 56.1 59.2 52.8 8.5
TGR-Token 25.4 33.2 35.6 31.1 29.1 37.0 39.5 34.8 41.2 53.0 55.8 49.3 7.5
TGR-Latent 27.8 37.4 40.1 34.8 32.8 41.8 44.9 39.5 43.8 56.8 60.1 52.8 7.0
Table 2: Code generation results on Qwen3-8B. TGR-Latent achieves the highest AUC on all three benchmarks and leads at high $k$, where latent-space diversity enables exploration of multiple valid coding solutions.
Method HumanEval BigCodeBench LiveCodeBench Tokens
@1 @32 @128 AUC @1 @32 @128 AUC @1 @32 @128 AUC Avg. ($\times 10^{3}$)
Base Model 45.7 67.8 71.2 60.3 34.1 48.5 51.3 43.8 26.8 36.8 38.9 33.5 1.6
PSampling 52.8 74.5 77.8 67.2 40.8 54.2 57.2 49.9 32.4 42.1 44.1 38.8 3.6
LCoT-GRPO 55.0 76.8 79.5 69.5 43.5 57.5 60.2 52.8 35.5 44.5 46.8 41.5 6.0
LCoT-SimKO 56.2 79.2 82.0 71.8 44.8 60.1 62.8 55.0 36.5 46.5 48.8 43.2 6.0
Dele-GRPO 56.4 79.8 82.5 71.2 44.8 60.5 63.3 55.1 36.4 46.8 49.1 43.4 5.9
Dele-SimKO 57.1 80.5 83.5 73.2 45.5 61.3 64.1 56.0 37.2 47.5 49.8 44.1 5.8
TGR-Token 55.3 78.5 81.8 70.8 43.5 58.8 61.4 53.6 35.1 45.5 47.8 42.0 4.4
TGR-Latent 56.1 81.1 84.2 73.5 45.6 62.1 64.8 56.3 36.8 48.1 50.2 44.4 4.6
4 Experiments
4.1 Experimental Setup
Models
We use Qwen3 (Yang et al., 2025) as the primary testbed and report main results for the 8B variant; additional scales, 1.7B, 4B and 14B, are deferred to Appendix LABEL:app:additional_results. To study the interaction with training stages, we also evaluate Phi-4-reasoning (SFT Pre-Training) and Phi-4-reasoning-plus (with RL Post-Training) (Abdin et al., 2025).
Baselines
We compare against (i) training-free test-time compute baselines: the base model and Power Sampling (PSampling) (Karan and Du, 2025); (ii) RL-tuned approaches that internalize long-CoT behaviors, including LCoT(GRPO) (Guo et al., 2025) and LCoT(SimKO) (Peng et al., 2025); (iii) RL with structured state transitions, instantiated by Delethink (Aghajohari et al., 2025) combined with GRPO or SimKO; and (iv) token-space control ablations, which mirrors strong test-time strategies like Self-Consistency (Wang et al., 2022), ToT (Yao et al., 2023) or GoT (Besta et al., 2024) that operate in token-space under the same chunked interface.
Benchmarks
For mathematics, we use MATH500 (Hendrycks et al., 2021), AIME2025 (MAA, 2025), OmniMath (Gao et al., 2024), and OlympiadBench (He et al., 2024). For code generation, we use HumanEval (Chen, 2021), BigCodeBench (Zhuo et al., 2024), and LiveCodeBench (Jain et al., 2024).
Unified inference budget and metrics.
We report Pass@$k$ for $k\in\mathcal{K}$ and the normalized area under the Pass@$k$ curve, namely AUC as robustness metrics, along with Avg. Tokens, defined as the average end-to-end tokens per problem required to produce $k$ final trajectories (Appendix LABEL:app:exp_grid).
We define the normalized AUC on a log scale as
$\text{AUC}\triangleq$ $\frac{100}{\log_{2}k_{M}-\log_{2}k_{0}}\sum_{i=0}^{M-1}\big(\log_{2}k_{i+1}-\log_{2}k_{i}\big)$ (10)
$\cdot\frac{\text{Pass}@k_{i+1}+\text{Pass}@k_{i}}{2},$
where $\mathcal{K}=\{k_{0}<\dots<k_{M}\}$. For code benchmarks, Pass@$k$ is computed using the unbiased estimator of Chen (2021). For math benchmarks, Pass@$k$ is the empirical success rate that at least one of $k$ samples is correct. We use temperature $0.6$ and the default TGR sampling with $K=8$ candidate anchors per chunk boundary, scored by rapid rollouts ($s=32$ on math tasks and $s=64$ on code tasks); full hyperparameters are listed in Appendix LABEL:app:hyperparams. We set an upper bound of $L=24$ chunks with maximum chunk length $S=512$, while actual generation terminates early by standard stopping criteria; all reported costs are based on realized token consumption.
4.2 Main Results
We evaluate TGR on math and code benchmarks under matched inference budgets. TGR-Latent consistently improves robust trajectory coverage while remaining token-efficient, indicating higher effective sample efficiency than both training-free sampling and training-heavy RL baselines. The comparison with TGR-Token isolates the benefit of latent-space search beyond discrete token-level control.
Mathematical Reasoning
Table 1 reports results on Qwen3-8B. TGR-Latent consistently improves robust coverage, with gains scaling at medium-to-large sampling budgets. This indicates effective conversion of samples into distinct solution trajectories. Moreover, TGR attains these gains with token costs competitive with training-heavy RL baselines.
Code Generation
Table 2 summarizes code generation results. TGR-Latent improves AUC consistently while maintaining comparable token cost. The advantage is pronounced on coding tasks, which often admit multiple correct solutions. Substantial gains at higher $k$ values are achieved by TGR through effective exploration of diverse coding trajectories in latent space.
Ablation of Geometric Components.
Table 3 isolates each score component. Removing look-ahead value $V_{\mathrm{fore}}$ yields the largest degradation, confirming that look-ahead guidance from limited-horizon look-ahead rollouts is the primary driver. Removing uniformity $P_{\mathrm{uni}}$ substantially reduces AUC, indicating that repulsion is essential for candidate pool diversity. Bumpiness $P_{\mathrm{bum}}$ stabilizes multi-chunk coherence. The gap between TGR-Latent and TGR-Token suggests that improvements stem from latent-space search, not merely chunking.
Table 3: Component ablation on Qwen3-8B. Lightweight foresight is essential for performance, while bumpiness and uniformity regularizers provide complementary gains. Latent-space control consistently outperforms its discrete token-level counterpart.
Method MATH500 AIME25 Avg. AUC $\Delta$
TGR-Latent 71.5 34.8 53.2 -
- w/o $P_{\mathrm{uni}}$ 67.2 31.0 49.1 - 7.6%
- w/o $P_{\mathrm{bum}}$ 69.1 31.7 50.9 - 4.2%
- w/o $V_{\mathrm{fore}}$ 61.3 26.1 43.7 - 17.8%
- random anchor 55.1 21.5 38.3 - 27.9%
TGR-Token 66.4 31.1 48.8 - 8.3%
Figure 3: TGR dominates the inference efficiency frontier. Left: Pass@$k$ curves on MATH500 reveal that TGR-Latent sustains marginal gains beyond $k=32$ where baselines plateau. Middle & Right: On the cost–robustness plane, TGR-Latent occupies the upper-left corner, achieving the highest AUC at moderate token cost on both math and code benchmarks.
4.3 In-Depth Analysis of TGR on Qwen3-8B
To understand TGR’s geometric design and clarify the source of its cost–robustness advantage, we interpret how soft geometric scoring balances foresight, smoothness, and diversity to sustain effective coverage under budget constraints.
Inference scaling on MATH500.
Figure 3 (left) plots Pass@$k$ on MATH500. The critical signal is the marginal gain as $k$ increases. TGR-Latent sustains larger improvements at medium-to-large $k$, consistent with higher candidate independence. The repulsive term minimizes overlap, ensuring additional samples translate into new trajectory modes. Value-based look-ahead sharpens boundary selection, converting rollout compute into sustained gains.
Cost–robustness frontier.
Figure 3 (middle, right) summarizes the cost–robustness trade-off. Methods relying on training-heavy policies can incur substantial token cost without proportional coverage increase, while sampling-based approaches saturate early. TGR-Latent makes the most of its compute budget, allocating it to lightweight look-ahead and pruning low-yield exploration. This supports the conclusion that soft geometric scoring improves the conversion rate from inference compute to trajectory coverage.
Mode collapse and search dynamics.
To diagnose diversity collapse, we analyze the geometry of within-step candidate anchors over time. As shown in Figure 4 (left), RL-tuned baselines concentrate probability mass in a narrow latent region, yielding highly correlated candidates and reduced coverage. In contrast, TGR maintains a dispersed anchor distribution, reflecting the effect of geometric repulsion in sustaining effective parallelism during scoring. This qualitative observation is quantified in Table LABEL:tab:latent_stats, where removing $P_{\mathrm{uni}}$ collapses the candidate set toward a narrow cone, while the full model preserves meaningful separation.
Hyperparameter sensitivity.
Figure 4 (right) reports sensitivity to rollout depth $s$ and the candidate budget $K$, defined as the number of candidate anchors scored per boundary. Increasing either parameter improves AUC but with diminishing returns, while overhead grows approximately linearly. Consistent with the design goal of budget-aware inference-time search, where moderate look-ahead is sufficient to rank anchors reliably, and soft geometric penalties maintain diversity without excessive sampling, the default configuration lies near a stable operating region.
Figure 4: Left: Latent-space mode diversity. RL-tuned baselines collapse into a unimodal cone, while TGR preserves a well-dispersed distribution, capturing a fuller range of valid reasoning paths. Right: Hyperparameter robustness. AUC increases with rollout depth $s$ and beam width $K$, but with diminishing returns.
4.4 Interpreting TGR Representations
Beyond performance gains, we seek to interpret how TGR’s geometric design shapes its internal representations and search behavior, shedding light on the mechanisms underlying robust coverage.
Balancing Value, Bumpiness, and Uniformity in TGR.
Soft geometric scoring manages a three-way trade-off among look-ahead value, cross-chunk coherence, and exploration. Greedy decoding over-optimizes near-horizon likelihood; aggressive sampling sacrifices coherence. TGR makes this trade-off explicit via a single objective combining rollout value with bumpiness and uniformity regularizers. Regularization weights $(\lambda_{b},\lambda_{u})$ control coherence and coverage. Empirically, removing $V_{\mathrm{fore}}$ degrades performance most, while removing $P_{\mathrm{uni}}$ primarily reduces AUC, indicating diminished parallelism.
Effective Independence Explains Stronger Gains in AUC than in Pass@1.
AUC aggregates performance across the sampling scale, proxying the effective number of distinct trajectories. When candidates are correlated, increasing $k$ yields diminishing returns. The uniformity term $P_{\mathrm{uni}}$ reduces overlap, ensuring additional compute translates into new solution modes. Thus, TGR’s advantage manifests primarily as higher AUC rather than Pass@$1$. Furthermore, Table LABEL:tab:latent_stats quantifies latent statistics showing that removing $P_{\mathrm{uni}}$ collapses effective candidate-set size $N_{\mathrm{eff}}$.
Geometry–robustness diagnostics.
We find that robustness exhibits a consistent geometric signature, with AUC peaking at an intermediate mean geodesic curvature $\kappa$ that is neither overly rigid nor overly erratic and increasing with the effective candidate-set size proxy $N_{\mathrm{eff}}$. This pattern supports the claim that TGR’s advantage stems from maintaining genuinely distinct search paths rather than redundant variations, as illustrated in Figure LABEL:fig:fa_curv_neff).
5 Discussion
5.1 Where Geometric Constraints Reside
Recent work interprets Transformer representations as trajectories on a latent manifold, with training objectives shaping its underlying geometry (Aggarwal et al., 2025; Di Sipio et al., 2025; Zhu et al., 2025; Xie et al., 2025). Building on this view, TGR operationalizes geometric structure at inference time through a manifold-informed scoring function over chunk-level candidates, enabling trajectory search without retraining. The key distinction lies in how geometric structure is instantiated. RL internalizes it into model weights, yielding strong single-trajectory priors but concentrating probability mass and limiting controllable exploration (Yue et al., 2025; Srivastava and Aggarwal, 2025). Training-time architectural regularization embeds it into the model parameterization, yet requires modifications to the training pipeline. In contrast, TGR externalizes geometric guidance as an inference-time control layer, offering tunable knobs that directly mediate the trade-off among look-ahead value, coherence, and diversity. This explicit control enhances robustness in long-horizon reasoning by increasing the effective independence of explored trajectories, naturally yielding higher AUC at medium-to-large sampling scales.
5.2 Training Stage Determines Inference-Time Controllability
This framing clarifies a key constraint on inference-time controllability, in which score guidance yields larger gains on SFT models than on heavily RL-optimized ones (Figure 5). Once training concentrates the trajectory distribution, degrees of freedom for inference-time search shrink; improvements manifest primarily as stabilizing coverage rather than shifting single-sample accuracy. Conversely, when training does not fully shape long-horizon behavior, explicit inference-time geometry provides a direct mechanism for allocating compute to high-yield branches while maintaining effective parallelism.
Figure 5: Training stage modulates inference-time controllability. TGR-Latent yields substantially larger gains on the SFT model (top), while improvement narrows after RL optimization (bottom), suggesting that inference-time search benefits models whose trajectory distribution retains residual flexibility.
5.3 Limitations and Future Directions
TGR trades training cost for inference overhead, impacting latency-sensitive serving. Low-overhead rollouts serve as efficient proxies for downstream utility but may be less informative under delayed credit; adaptive budgeting or search at high-uncertainty boundaries offer mitigations.
Agentic workflows offer significant potential, yet failures frequently arise prior to tool execution due to unstable or ambiguous intent. A low-overhead latent foresight step guided by soft geometric principles could provide a reliable intent anchor prior to tool invocation, offering a potential mechanism to stabilize downstream planning.
6 Conclusion
We introduced TGR, a training-free inference-time framework that steers long-horizon reasoning via manifold-informed latent foresight search. By scoring chunk-level anchors with a soft geometric objective combining look-ahead value, coherence, and diversity penalties, TGR matches or exceeds RL-tuned baselines on challenging math and code benchmarks with negligible overhead. Decoupling lightweight boundary rollouts and geometric preferences from model training reveals that robust coverage need not be amortized through training but can instead be induced on the fly via structured latent-space search.
Impact Statement
This paper studies a training-free, manifold-informed latent foresight search framework for long-context reasoning. If effective, it may improve robustness under fixed or modestly increased compute budgets, benefiting applications such as programming assistance. As with other capability-improving methods, it could also enable misuse (e.g., harmful or misleading content) and its impact depends on deployment context. We recommend pairing such techniques with safeguards and careful evaluation before deployment.